the perimeter of a square and circle are the same. what is the ratio between the area of the circle to the area of the square?
Answer by rothauserc(4718)

Perimeter of square = 4s, where s is the length of a side
:
Perimeter of a circle = 2*pi*r, where pi=(22/7) and r is the radius, then
:
4s = 2*pi*r
:
s = (pi*r) / 2
:
Area of square is s^2 and Area of circle is pi*r^2, then
:
(pi*r^2) / ((pi*r /2))^2 = (pi*r^2) / ((pi^2 * r^2)/4) =
:
((pi*r^2) * 4) / (pi^2 * r^2) = 4 / pi = (4 * 7) / 22 = 28 / 22 = 14 / 11
:
****************************************************************
Ratio of Area of the circle to the Area of the square is 4 / pi
which is 14 / 11
****************************************************************
:
***************************************************************
: